# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ A CTMC admits a Poincaré constant $C_{P}$ if for all test function $f\in\ell_{2,\nu}$, we have
(17) $\mathcal{E}(f,f)=\langle f,-\mathcal{L}f\rangle_{\nu}=-\sum_{x,y\in\mathcal{S}}\nu(x)\mathcal{L}(x,y)f(x)f(y).$
It is known that Poincaré inequalities immediately imply exponential ergodicity of the CTMC (Bakry et al., 2013). We formally state it in the following proposition for completeness.
@@
(18) $\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu)$
@@
The proof of Proposition 6 is well-known and can be found in (Bakry et al., 2013) and in Appendix A.1. Proposition 6 shows that Poincaré inequality with constant $C_{P}$ implies exponential convergence in the chi-square distance with mixing rate $1/C_{P}$. Thus, for the rest of the section, our focus is on establishing the Poincaré constant.
Building on the recent developments on Markov chain mixing (Taghvaei and Mehta, 2022; Andrieu et al., 2022), we now present the Lyapunov-Poincaré methodology below, which has two main ingredients. The first one is a Foster-Lyapunov like assumption:
@@
(19) $\mathcal{L}V(q)\leq-\gamma V(q)+b(q)\,\forall q\in\mathcal{S}.$
@@
(20) $\mathcal{L}V(q)\leq-\gamma V(q)+B1_{\{x^{*}\}}\,\forall q\in\mathcal{S}.$
@@
While a similar convergence result for TV distance was previously established (Lund et al., 1996), this result allows us to establish Chi-square convergence, and we present the proof in Appendix A.2.3. Combining the above result with Proposition 6, one can obtain the Chi-square convergence as in (18). More generally, the finite set $K$ may not be a singleton for most of the queueing systems (including the $M/M/n$ queue), and thus, the Foster-Lyapunov assumption alone fails to provide a Poincaré inequality since we do not know how the system would mix inside $K$. So we need to establish a local mixing result inside $K$, which can be done using a local Poincaré inequality. Such a result is formalized in the following assumption.
@@
(21) $\operatorname{Var}_{\tau}(f)=\left\lVert f-\mathbb{E}_{\tau}[f]\right\rVert_{2,\tau}^{2}\leq\frac{C_{b}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}.$
@@
(22) $\operatorname{Var}_{\nu}(f)\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$
@@
The Stitching Theorem (Theorem 1) is applicable for irreducible, aperiodic, positive recurrent and reversible CTMCs. For the purpose of analyzing queueing systems, we state the Stitching Theorem for the countable state space Markov chains, but this theorem can be easily generalized to the continuous state space (Taghvaei and Mehta, 2022) and can potentially be applied to other areas such as Learning and Sampling (Durmus and Moulines, 2014; Raginsky et al., 2017; Vempala and Wibisono, 2022). The proof of Theorem 1 can be found in Appendix A.2.1. For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, Assumption 4 gives the local Poincaré inequality
(23) $\operatorname{Var}_{\nu_{K}}(f)\leq\frac{C_{L}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu},$
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$. And so, this allows us to recover the result of (Taghvaei and Mehta, 2022) as stated below.
@@
A version of Corollary 2 is known in (Taghvaei and Mehta, 2022; Rosenthal, 1995) for reversible Markov chains and (Dominique et al., 2008; Raginsky et al., 2017) for Markov processes. The readers can find the proof of Corollary 2 in Appendix A.2.2, which is by applying the result of Theorem 1 for $b=B1_{K}$. The constant $C_{L}$ is called the local Poincaré constant w.r.t. the finite set $K$ of the CTMC. Intuitively, the CTMC mixes fast outside the finite set $K$ due to the negative drift, in addition to the rapid mixing within $K$ ensured by the “local” Poincaré inequality. In many applications, establishing a tight characterization of the local mixing behavior is the key to obtaining the tight mixing rate, and previously, the local Poincaré constant is predominantly obtained using the minorization condition (Rosenthal, 1995; Down et al., 1995; Taghvaei and Mehta, 2022). However, it is known that the approach using minorization condition often yields sub-optimal bounds, and the minorization condition itself is usually hard to obtain as well (Qin and Hobert, 2020; Anderson et al., 2020). Moreover, whenever the bounded set is not a singleton, there will be a ”stitching” error when combining the mixing behavior inside the finite set with the mixing behavior outside the finite set together. Therefore, our Stitching Theorem (Theorem 1) is a necessary generalization to achieve a tight characterization of the local mixing behavior inside the finite set, whereas a naive application of Corollary 2 would not yield such a result. For a more detailed explanation, we refer the readers to the discussion around Equation (29).
@@
Comparing Corollary 2 with Proposition 7, one can see that when we try to ”stitch” negative drift outside of the finite set $K$ with the local mixing condition behavior inside the finite set, we incur a ”stitching error” $BC_{L}$ which hurts our Poincaré constant. Ideally, to get a tight mixing bound, we would like to make this stitching error term to be as small as possible or vanishes to 00 as $n$ goes to infinity. And so, it is crucial to choose the right Lyapunov function so that we get that right negative drift rate $\gamma$ while minimizing $B$, as well as developing methods to get tight local mixing constant $C_{L}$. We shall elaborate more on these steps in the following Section 4.
4. Proof sketch of mixing results
@@
Recall that the $M/M/n$ queue experiences a phase transition at the Halfin-Whitt regime, i.e. $\alpha=\frac{1}{2}$. And so, to prove our main results, we will take two different approaches: a Lyapunov-Poincaré approach for the Super-Halfin-Whitt regime and a Lyapunov coupling approach for the Sub-Halfin-Whitt regime. The intuitions behind our approaches and the reasons why these regimes require two different approaches will become apparent as we go through the subsections below.
4.1. Regime 1: Super-Halfin-Whitt Regime
@@
(24) $\mathcal{L}V\leq-\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}V+b1_{\{0\}}$
where $b$ is some universal constant. The rate of $\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}$ in the above equation matches previously established mixing results (Van Doorn, 1985; Robert, 2003) for the $M/M/1$ queue suggesting that $V$ is indeed the correct Lyapunov function. As seen in (24), we have a negative drift everywhere except the singleton $\{0\}$, so we simply apply Proposition 7 to obtain the Poincaré constant, which gives us the desired mixing rate.
@@
(25) $V(q)=e^{\theta\left[q-(n-1)\right]^{+}+\theta\left[(n-1)-q\right]^{+}}$
with appropriately chosen $\theta\geq 0$ to accurately capture the negative drift above $\lambda$ and the positive drift below $\lambda$. While $\lambda$ is the center of the drift, we consider the distance of the queue length to $n-1$ in the exponent because of algebraic convenience in the proof. From this choice of Lyapunov function, we obtain the following drift lemma for the Super-Halfin-Whitt regime.
@@
(26) $\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}\,\forall q\in\mathbb{Z}_{\geq 0}$
where
(27) $K=\begin{cases}\{0,1,...,n-1\}&\text{ if }1\leq n\leq 7\\ \{\lceil 2\lambda_{n}\rceil-n,...,n-1\}&\text{ if }n>7.\end{cases}$
Furthermore, we have $b(q)\leq L\,\forall q\in K$ for some constant $L>0$ and $2\leq n\leq 7$ and for $n>7$, we have
(28) $b(q)=\begin{cases}2e^{2}n^{1-2\alpha}\,\,&\forall q\in K-\{n-1\}\\ 3n^{1-\alpha}&\text{if }q=n-1\\ 0\,\,&\forall q\not\in K.\end{cases}$
The above lemma completes Step 1 of our methodology by establishing a negative drift with rate $\gamma_{n}$ outside the finite set $K$. Unlike (Taghvaei and Mehta, 2022), the function $b$ in our drift condition (26) is not a constant function, which allows a more refined analysis and precipitates the use of our Stitching Theorem 1. To understand why the chosen Lyapunov function works for the $M/M/n$ queue, consider the following ODE that is a fluid analog of $M/M/n$ queue: $\dot{q}=\lambda_{n}-\min\{q,n\}$. When $q\geq\lambda_{n}+n^{1-\alpha}$ and $\theta_{1}\approx n^{-\alpha}$, we have $\dot{q}=\lambda_{n}-n=-n^{1-\alpha}$ which gives $\dot{V}\approx-n^{1-2\alpha}V\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$, which is the desired negative drift. Similarly, when $q\leq\lambda_{n}-n^{1-\alpha}$, we also have $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$. This allows us to establish a negative drift with rate $-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ outside of a set $K$ with cardinality at most $\lceil 2n^{1-\alpha}-1\rceil$.
@@
However, note that from (28), if we choose $B=\max_{q\in\mathbb{Z}_{\geq 0}}b(q)$ and then naively apply Theorem 1 in (Taghvaei and Mehta, 2022) (also Corollary 2 in our work), we would get a Poincaré constant
(29) $C_{P}=\frac{1+O\left(n^{2-3\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$
@@
(Weighted Poincaré for $M/M/n$ in the Super-Halfin-Whitt regime) Under Assumptions 1, 2, let $b$ be given by (28), $n\geq 65$ and $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$. In addition, denote $\nu_{n}$ as the stationary distribution , we have Assumption 4 holds with constant $C_{b}=\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}$ holds for $\alpha\in[1/2,1)$ and for all test function $f\in\ell_{2,\nu_{n}}$ and for some positive constants $\Delta_{1},\Delta_{2}$.
Note that we assume $n\geq 65$ only to obtain reasonable universal constants in the weighted-Poincaré constant. Obtaining an explicit bound for all $n\geq 1$ is possible, but doing so would result in unreasonably large universal constants. One can further improve such universal constants by increasing the threshold depending on the applications. The proof and the precise characterization of the constant $C_{b}$ is reserved in Subsection 5.1, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 1 to obtain the results in Equation (6) and Equation (7) in Theorem 3 which gives us a convergence bound for $\alpha\in(1/2,\infty)$ that matches the limiting behavior at the asymptotics. The full detail of the proof is deferred to Appendix B.1.4.
@@
Remark: Previously, several works have attempted different methods to establish the local Poincaré inequality. Previous works (Taghvaei and Mehta, 2022; Rosenthal, 1995) rely on the minorization assumption to establish some form of local mixing inside $K$, i.e. a local Poincaré inequality as in (23). However, such an approach does not directly work in our setting since the assumption requires every state in the state space to be within reach of any other state, which clearly does not hold for a birth-and-death process. Such an assumption can be satisfied if one looks at an $m-$step transition matrix for a large enough $m>0$. However, such an approach usually results in a poor minorization constant. In addition, a few other works have obtained local Poincaré inequalities using the diameter of the finite set, such as (Raginsky et al., 2017; Divol et al., 2024), but generally, the local Poincaré constant bounds in these works are loose. Thus, as our goal is to obtain a tight characterization of the mixing rate, we take a different approach and propose a local version of the canonical path method that exploits the birth-and-death structure of the transition matrix. In fact, the non-local version has been used to bound the mixing time of random walks on different graph topologies (Berestycki, 2020).
@@
The mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ in the regime $\alpha\geq 1$ is tight and matches the result in (Zeifman, 1991). In addition, the mixing rate for the regime $\alpha\in(1/2,1)$ is asymptotically tight, i.e., it matches with $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ for large $n$. For $\alpha=1/2$, the Poincaré constant inside the finite set $K$ does not vanish. We show that it is bounded above by a constant implying that the mixing rate is bounded below by a universal constant.
4.2. Regime 2: Sub-Halfin-Whitt Regime

### v4

@@ A CTMC admits a Poincaré constant $C_{P}$ if for all test function $f\in\ell_{2,\nu}$, we have
$\mathcal{E}(f,f)=\langle f,-\mathcal{L}f\rangle_{\nu}=-\sum_{x,y\in\mathcal{S}}\nu(x)\mathcal{L}(x,y)f(x)f(y).$ (21)
It is known that Poincaré inequalities immediately imply exponential ergodicity of the CTMC [3]. We formally state it in the following proposition for completeness.
@@
$\chi^{2}\left(\pi_{t},\nu\right)\leq e^{-\frac{2t}{C_{P}}}\chi^{2}(\pi_{0},\nu)$ (22)
@@
The proof of Proposition 6 is well-known and can be found in [3] and in Appendix 8.1. Proposition 6 shows that Poincaré inequality with constant $C_{P}$ implies exponential convergence in the chi-square distance with mixing rate $1/C_{P}$. Thus, for the rest of the section, our focus is on establishing the Poincaré constant.
Building on the recent developments on Markov chain mixing [71, 2], we now present the Lyapunov-Poincaré methodology below, which has two main ingredients. The first one is a Foster-Lyapunov like assumption:
@@
$\mathcal{L}V(q)\leq-\gamma V(q)+b(q)\,\forall q\in\mathcal{S}.$ (23)
@@
$\mathcal{L}V(q)\leq-\gamma V(q)+B1_{\{x^{*}\}}\,\forall q\in\mathcal{S}.$ (24)
@@
While a similar convergence result for TV distance was previously established [44], this result allows us to establish Chi-square convergence, and we present the proof in Appendix 8.2.3. Combining the above result with Proposition 6, one can obtain the Chi-square convergence as in (22). More generally, the finite set $K$ may not be a singleton for most of the queueing systems (including the $M/M/n$ queue), and thus, the Foster-Lyapunov assumption alone fails to provide a Poincaré inequality since we do not know how the system would mix inside $K$. So we need to establish a local mixing result inside $K$, which can be done using a local Poincaré inequality. Such a result is formalized in the following assumption.
@@
$\operatorname{Var}_{\tau}(f)=\left\lVert f-\mathbb{E}_{\tau}[f]\right\rVert_{2,\tau}^{2}\leq\frac{C_{b}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu}.$ (25)
@@
$\operatorname{Var}_{\nu}(f)\leq\frac{1+\left(\sum_{q\in\mathcal{S}}b(q)\nu_{K}(q)\right)C_{b}}{\gamma}\langle f,-\mathcal{L}f\rangle_{\nu}\,\forall f\in\ell_{2,\nu},$ (26)
@@
The Stitching Theorem (Theorem 2) is applicable for irreducible, aperiodic, positive recurrent and reversible CTMCs. Our Stitching Theorem is a generalization of Theorem 1 in [71] for countable state space CTMCs and this generalization is crucial in getting the tight mixing rate in the Super-Halfin-Whitt regime all the way to $\alpha>1/2$ instead of $\alpha>2/3$. We defer this discussion to (32) and the proof of Theorem 2 to Appendix 8.2.1. For the special case of $b=B1_{K}$ for some finite $K\subseteq S$ and $B>0$, Assumption 4 gives the local Poincaré inequality
$\operatorname{Var}_{\nu_{K}}(f)\leq\frac{C_{L}}{\nu(K)}\langle f,-\mathcal{L}f\rangle_{\nu},$ (27)
where $\nu_{K}(x)=\frac{\nu(x)}{\sum_{x\in K}\nu(x)}\forall x\in K$. And so, this allows us to recover the result of [71] as stated below.
@@
A version of Corollary 7 is known in [71, 68] for reversible Markov chains and [21, 66] for Markov processes. The readers can find the proof of Corollary 7 in Appendix 8.2.2, which is by applying the result of Theorem 2 for $b=B1_{K}$. The constant $C_{L}$ is called the local Poincaré constant w.r.t. the finite set $K$ of the CTMC. Intuitively, the CTMC mixes fast outside the finite set $K$ due to the negative drift, in addition to the rapid mixing within $K$ ensured by the “local” Poincaré inequality. In many applications, establishing a tight characterization of the local mixing behavior is the key to obtaining the tight mixing rate, and previously, the local Poincaré constant is predominantly obtained using the minorization condition [68, 23, 71]. However, it is known that the approach using minorization condition often yields sub-optimal bounds, and the minorization condition itself is usually hard to obtain as well [63, 1]. Moreover, whenever the bounded set is not a singleton, there will be a ”stitching” error when combining the mixing behavior inside the finite set with the mixing behavior outside the finite set together. Therefore, our Stitching Theorem (Theorem 2) is a necessary generalization to achieve a tight characterization of the local mixing behavior inside the finite set, whereas a naive application of Corollary 7 would not yield such a result. For a more detailed explanation, we refer the readers to the discussion around Equation (32).
@@
Comparing Corollary 7 with Proposition 7, one can see that when we try to ”stitch” negative drift outside of the finite set $K$ with the local mixing condition behavior inside the finite set, we incur a ”stitching error” term $BC_{L}$ which hurts our Poincaré constant. Ideally, to get a tight mixing bound, we would like to make this stitching error term to be as small as possible or vanishes to 0 as $n$ goes to infinity. And so, it is crucial to choose the right Lyapunov function so that we get that right negative drift rate $\gamma$ while minimizing $B$, as well as developing methods to get tight local mixing constant $C_{L}$. We shall elaborate more on these steps in the following Section 4.
4 Proof sketch of mixing results
@@
Recall that the $M/M/n$ queue experiences a phase transition at the Halfin-Whitt regime, i.e. $\alpha=\frac{1}{2}$ [32]. And so, to prove our main results, we will take different approaches and construct different Lyapunov functions for each of these regimes. The intuitions behind our approaches and the reasons why these regimes require two different approaches will become apparent as we go through the subsections below.
Figure 1: Visualization of chosen Lyapunov functions for the Super-Halfin-Whitt regime, Sub-Halfin-Whitt regime and the Mean Field regime. Figure 1(a) corresponds to the piecewise exponential Lyapunov function in the Super-Halfin-Whitt regime. Figure 1(b) corresponds to the linear and exponential Lyapunov function in the Sub-Halfin-Whitt regime. Figure 1(c) corresponds to the constant and exponential Lyapunov function in the Mean-Field regime.
4.1 Regime 1: Super-Halfin-Whitt Regime
@@
$\mathcal{L}V\leq-\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}V+b1_{\{0\}}$ (28)
where $b$ is some universal constant. The rate of $\left(\sqrt{n}-\sqrt{\lambda}\right)^{2}$ in the above equation matches previously established mixing results [73, 67] for the $M/M/1$ queue suggesting that $V$ is indeed the correct Lyapunov function. As seen in (28), we have a negative drift everywhere except the singleton $\{0\}$, so we simply apply Proposition 7 to obtain the Poincaré constant, which gives us the desired mixing rate.
@@
$V(q)=e^{\theta\left[q-(n-1)\right]^{+}+\theta\left[(n-1)-q\right]^{+}}$ (29)
with appropriately chosen $\theta\geq 0$ to accurately capture the negative drift above $\lambda$ and the positive drift below $\lambda$. While $\lambda$ is the center of the drift, we consider the distance of the queue length to n1n-1 in the exponent because of algebraic convenience in the proof. From this choice of Lyapunov function, we obtain the following drift lemma for the Super-Halfin-Whitt regime.
@@
$\mathcal{L}V(q)\leq-\gamma_{n}V(q)+b(q)1_{K}\,\forall q\in\mathbb{Z}_{\geq 0}$ (30)
where $K=\{[\lfloor 2\lambda_{n}\rfloor-n]^{+},...,n-1\}$. Furthermore, we have $b(q)\leq L\,\forall q\in K$ for some constant $L>0$ and $2\leq n\leq 7$ and for $n>7$, we have
$b(q)=\begin{cases}\frac{n^{1-2\alpha}}{1-n^{-\alpha}}e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}\,\,&\forall q\in K-\{n-1\}\\ n^{1-\alpha}\bigg{(}1+\frac{1}{4(n^{\alpha}-1)}\bigg{)}&\text{if }q=n-1\\ 0\,\,&\forall q\not\in K.\end{cases}$ (31)
The above lemma completes Step 1 of our methodology by establishing a negative drift with rate $\gamma_{n}$ outside the finite set $K$, the detail proof of this Lemma is deferred to Subsection 5.1. Unlike [71], the function $b$ in our drift condition (30) is not a constant function, which allows a more refined analysis and precipitates the use of our Stitching Theorem 2. To understand why the chosen Lyapunov function works for the $M/M/n$ queue, consider the following ODE that is a fluid analog of $M/M/n$ queue: $\dot{q}=\lambda_{n}-\min\{q,n\}$. When $q\geq\lambda_{n}+n^{1-\alpha}$ and $\theta_{1}\approx n^{-\alpha}$, we have $\dot{q}=\lambda_{n}-n=-n^{1-\alpha}$ which gives $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V=-\Theta(n^{1-2\alpha})V$, which is the desired negative drift. Similarly, when $q\leq\lambda_{n}-n^{1-\alpha}$, we also have $\dot{V}\leq-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}V$. This allows us to establish a negative drift with rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ outside of a set $K$ with cardinality at most $\lceil 2n^{1-\alpha}-1\rceil$.
@@
However, note that from (31), if we choose $B=\max_{q\in\mathbb{Z}_{\geq 0}}b(q)$ and then naively apply Theorem 1 in [71] (also Corollary 7 in our work), we get $B=b(n-1)=O(n^{1-\alpha})$. Combining with the fact that $C_{L}=\Theta(n^{1-2\alpha})$, we would get the Poincaré constant
$C_{P}=\frac{1+O\left(n^{2-3\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$ (32)
@@
(Weighted Poincaré for $M/M/n$ in the Super-Halfin-Whitt regime) Under Assumptions 1, 2, let $b$ be given by (31), and $K=\{\lfloor 2\lambda_{n}\rfloor-n,...,n-1\}$. In addition, denote $\nu_{n}$ as the stationary distribution, we have Assumption 4 holds with constant $C_{b}=\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}$ holds for $\alpha\in[1/2,1)$ and for all test function $f\in\ell_{2,\nu_{n}}$ and for some positive constants $\Delta_{1},\Delta_{2}$.
Since $\alpha>1/2$, we have $C_{b}\rightarrow 0$ as $n\rightarrow\infty$, which allows our finite-time mixing bound to match the limiting behavior at the asymptotic. The proof and the precise characterization of the constant $C_{b}$ is reserved in Appendix 5.2.3, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 2 to obtain the results in Equation (6) and Equation (7) in Theorem 1 and Proposition 2 as follows.
Finally, we bound the $\sum_{q\in K}b(q)\nu_{K}(q)$ term in Theorem 2, which is no more than a constant.
Claim 1.
Let $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$, we have $\sum_{q\in K}\tau(q)\leq g_{3}(n)$. Here, we have $g_{3}(n)\leq 3.48\,\forall n\geq n_{0}$ and $\lim_{n\rightarrow\infty}g_{3}(n)=1/2$ for $\alpha>1/2$ and $\lim_{n\rightarrow\infty}g_{3}(n)=e/2+1$ for $\alpha=1/2$.
With Lemma 1, Lemma 2 and Claim 1 (whose proof is in Appendix 9.1.2), we obtain the mixing rate of $M/M/n$ in the Super-Halfin-Whitt regime and $M/M/1$ as follows.
Proof of Theorem 1 for $\alpha\in(1/2,1)$ and Proposition 2:.
Case 1 ($\alpha\geq 1$ or $n=1$): From Proposition 7, Lemma 1 and observe that in this case, we have $K=\{n-1\}$, we have that the system admits the Poincaré constant $C_{P}=\frac{1}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$. And so, from Proposition 6, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n}).$ (33)
$C_{P}$ $=\frac{1+\left(\sum_{q\in K}b(q)\nu_{K}(q)\right)C_{b}}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}\leq\frac{1+3.48\times\left(384n^{2-4\alpha}+395.93n^{3-6\alpha}\right)}{(\sqrt{n}-\sqrt{\lambda_{n}})^{2}}$
Let $C_{n}=\frac{1}{1+3.48\times\left(384n^{2-4\alpha}+395.93n^{3-6\alpha}\right)}$, we have that $\lim_{n\rightarrow\infty}C_{n}=1$ since $\alpha>1/2$. From Proposition 6, we have
$\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})$ (34)
where $\lim_{n\rightarrow\infty}C_{n}=1$. ∎
@@
Remark: Previously, several works have attempted different methods to establish the local Poincaré inequality. Previous works [71, 68] rely on the minorization assumption to establish some form of local mixing inside $K$, i.e. a local Poincaré inequality as in (27). However, such an approach does not directly work in our setting since the assumption requires every state in the state space to be within reach of any other state, which clearly does not hold for a birth and death process. Such an assumption can be satisfied if one looks at an $m$-step transition matrix for a large enough $m>0$. However, such an approach usually results in a poor minorization constant. In addition, a few other works have obtained local Poincaré inequalities using the diameter of the finite set, such as [66, 20], but generally, the local Poincaré constant bounds in these works are loose. Thus, as our goal is to obtain a tight characterization of the mixing rate, we take a different approach and propose a local version of the canonical path method that exploits the birth and death structure of the transition matrix. In fact, the non-local version has been used to bound the mixing time of random walks on different graph topologies [6].
@@
In summary, the mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ in the regime $\alpha\geq 1$ is tight and matches the result in [86]. In addition, the mixing rate for the regime $\alpha\in(1/2,1)$ is asymptotically tight, i.e., it matches with $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ for large $n$. For $\alpha=1/2$, the Poincaré constant inside the finite set $K$ does not vanish. We show that it is bounded above by a constant implying that the mixing rate is bounded below by a universal constant.
4.2 Regime 2: Sub-Halfin-Whitt Regime